a researcher selects a sample of 32 participants who are assigned to participate in a study with one group. what are the degrees of freedom for this test?
Answer #1

Solution:

One sample test.

sample size n = 32

The degrees of freedom are the number of independent observations.

So , for one sample test,

Degrees of freedom = n - 1 = 32 - 1 = 31